Hello,
I need help finding the coordinates of the vertex of the graph of y=6x^2-14x.

Thank you

Answer by scott8148(5891)

the equation for the axis of symmetry (which passes through the vertex) is x=-b/2a

x=14/12 or x=7/6; this is the x value for the vertex

to find the y value, substitute into the original equation y=6(7/6)^2-14(7/6)